What are your preferred date(s) for proof re-test? As the beginning of the semester, would you have appreciated a “Web-based, artificially intelligent assessment and learning system,” for reviewing specific topics such as logs and partial fractions? If so, how much would you be willing to pay for this? On a scale of 1-10, is the class too fast/too slow?

too slow = 1, too fast = 10

On a scale of 1-10, is the algebra too fast/too slow?

too slow = 1, too fast = 10

How many hours per week do you spend on this course (include class attendance)? How many hours did you spend on the computer project. Other comments/suggestions: